1. Complete ratio analysis for the years 2010 through 2014, calculating the following ratios
    (2 marks each).

    Use the unaudited data for 2014. For any ratios that use an average of 2 years in the formula, you can assume that the 2010 ending balance approximates the 2010 average. (The ratios that use an average of 2 years in the formula include accounts receivable turnover, inventory turnover, return on assets, and return on equity.) You are not given the 2009 data to calculate the average for 2010, so you can assume that the average for 2010 is equal to the year-end balance. Show your calculations for each ratio or complete in Excel worksheet using formulas:
    1. Current ratio
    2. Receivables turnover
    3. Day’s sales in receivables
    4. Inventory turnover
    5. Inventory holding period
    6. Rate of return on net sales
    7. Rate of return on total assets
    8. Rate of return on stockholders’ equity
    9. Asset turnover
    10. EPS
    11. Debt ratio
    12. Times interest earned
  2. Provide a discussion of the company’s performance in terms of liquidity, profitability, and leverage. Be sure to discuss each ratio. In your discussion, identify the key indicators that showed the downfall of Forge Group. (6 marks)

Table 2. Comprehensive Income Statement (in thousands of Australian dollars) Unaudited June 30, 2010 June 30, 2011 June 30, 2012 June 30, 2013 January 31, 2014 $246,169 $421,595 $774,879 $1,054,100 $520,041 (711,430) Revenue Cost of sales Changes in inventories of finished goods and WIP Materials, plant, and contractor costs Employee benefits expense 9,696 (125,171) 79,194) (3,218) 15,000 (211,000) (157,191) 5,159) 516,867) (164,502) (16,292) (256,515) (21,361) Consulting fees Provision for impaiment losses Other expenses Other gains and losses (1,628) 7,132) (12.711) 21,033) 188 Results from Operating Activities S64,182 1,023 (716) 3,079 Finance income Finance costs Net finance income Share of profitAloss) of associates and jointly controlled entities 4,816) 2,123 (5,679) $90,109 27,190) $62,919 1,826 $64,745 $2,848 513 $56,753 17,920) Net Proit Before Tax $70,082 S(324,162) 2,301) (326,463) Income tax expense (10,915 $29,451 Net Profit After Tax Foreign exchange differences (net of tax) Total Comprehensive Income (1,946) $29,105

Table 3. Balance Sheet (in thousands of Australian dollars) June 30, 2010June 30, 2011 June 30,2012 June 30, 2013 January 31, 2014 Current Assets Cash and cash equivalents Short-term deposits Trade and other receivables Inventories and WIP Current tax assets Other assets Noncurrent assets classified as held for sale $51,921 $78,285 $51,091 72,500 196,884 11,331 $15,316 42,162 14,621 49,542 29,622 2.748 83,254 150,491 103,279 40,616 2,246 6,900 $117,850 1,574 2,487 1,560 (23,415) Total Current Assets $159,023 $331,316 $135,796 Noncurrent Assets Trade and other receivables Term deposits Property plant, and equipment Deferred tax assets Investments accounted for using equity method Intangibles Other assets 7,051 14,260 67,736 4,273 2,545 1,424 10,468 71,546 9,124 73,293 26,789 1,827 36,577 2,043 15,621 15,637 90,467 163,760 44,237 $162,087 144,108 $478,401 Total Noncurrent Assets 54,257 132,894 SA64,210 Total Assets

Current Liabilities Trade and other payables 52,968 72,845 267,169 219,56B 11.139 299,909 Current tax liabilities 3,970 Other liabilities 63,731 $363,640 Total Curent Liabilities $285,095 $234,677 Noncurrent Liabilities 9,246 17,453 Trade and other payables 1,517 4,901 14,547 1,067 Deferred tax liabilities Investments accounted for using the equity method 164 Other liabilities Total Noncurrent Liabilities $3,785 $29,981 $16,107 $52,184 Total Liabilities Net Assets Equity 415,824 (116,268) 68,711 88,818 124,426 315,076 250,784 213,426 5,430 5,430 62,919 1.471 103,606 213,426 86,169,014 Issued capital Profit reserve 42,839 44,294 1,034 49,505 912 81,080 124462 81,541,569 $5.46 1.221) 119,116 159,036 Retained earnings Total Equity Number of shares Share priCe (116,268) 86,169,014 70,699.,487 86,169,014
Answer #1

To calculate the following ratios-

(a) Current Ratio-

Current Asset/ Current Liabilities

Current Ratio for 2010 Current Ratio for 2014   

=117850/ 64926= 1.81 times = 135796/ 363640= .37 times

(b) Receivable turnover

= Net Credit Sales/ Average Accounts receivable

For 2010 For 2014

=246169 / 42612 = 5.77 times = 520041/ 93267= 5.57 times

  Average Accounts receivable for 2010- $ 42162

Average Accounts receivable for 2014- $ 83254 +103279/2 =93267

(c) Days sale in receivable

= Accounts receivable / Average sales per day

For 2010-

Accounts receivable= 42162

Average sales per day = 246169/365= 674.43

  Days sale in receivable = 42162/ 674.43 = 62.55

For 2014-

Accounts receivable= 73293

Average sales per day = 520041 /365= 1425

  Days sale in receivable= 73293/ 1425= 51.43

(d) Inventory Turnover Ratio

= Sales/ Average Inventory

For 2010-

246169/ 14621= 16.84

For 2014

=520041/ 93267 =5.57

(e) Rate of return on net sales

For 2010- For 2014-

Profit Margin/ Net sales *100

=29451/ 246169 *100 = 11.96 % = (326463)/ 520041 *100= 62.77%

(f) Rate of return on Total asset

  Profit Margin/ Total Asset  *100

  =29451/ 162087 *100 = 18.16 % = (326463)/ 299556 *100= 108.98%

(g) Rate of return on stock holder equity

  Net Income / stock holder equity *100

  =29451/ 93376 *100 = 31.54 % = (326463)/ 97158 *100= -336.01%

Average Equity for 2010 =93376

Average Equity for 2014= ( 213426-116268)/2 =97158

(h) Asset Turnover Ratio

= Revenue/ Net Asset

For 2010 For 2014  

=246169/ 162087= 1.51 =520041/ 299556= 1.73

(i) EPS

= Net Income- Preferred dividend/ Weighted Average equity shares

(j) Debt Ratio

For 2010 For 2014

= Total Debt/ Total Asset  

= 68711 /162087= .42 = 415824 /299556= 1.389

(b) Liquidity -

Company's liquidity is measured in terms of current ratio- It is a liquidity ratio that measures a company's ability to pay short-term and long-term obligations. As the current ratio is decreased company is facing liquidity issues.

Profitability

As the company is facing loss in 2014 - all the ratios have decreased. Company is having negative profitability ratios.

Leverage-

Company debt ratio is increased. The debt ratio is a financial ratio that measures the extent of a company's leverage. The lower the percentage, the less leverage a company is using and the stronger its equity position.

Hence, all this indicates that company is facing downfall.